# VISAGE: Video Instance Segmentation with Appearance-Guided Enhancement

### 08 Dec 2023 v1

@@ 4.3 Main Results
We compare our methods with state-of-the-art methodsg on four VIS benchmarks: YTVIS 2019/2021/2022 and OVIS. The results are reported in Tab. 1, Tab. 2 and Tab. 3.
Youtube-VIS 2019 & 2021.
As shown in Tab. 1, we compare our VISAGE with state-of-the-art methods. On the YTVIS 2019 benchmark, VISAGE demonstrates performance comparable to the highest-performing method [36]. Notably, both our VISAGE method and the best-performing method [36], surpass previous methods by a large margin. In the YTVIS 2021 benchmark, an improved version of YTVIS 2019, VISAGE outperforms other existing methods. We achieve competitive performance on both the YTVIS 2019 and 2021 datasets, by incorporating appearance-guided enhancement.
Youtube-VIS 2022.
The results for the YTVIS 2022 benchmark are presented in Tab. 2. This benchmark contains longer videos compared to previous benchmarks (YTVIS 2019/21). Our VISAGE exhibits slightly lower performance compared to GenVIS [16] and TCOVIS [23]. These methods utilize the entire history of previous information by propagating the queries. Nevertheless, our method employs only a fixed number of previous frames and shows comparable performance.
@@
VISAGE 34.8 59.8 35.3 16.2 39.4
Table 3: Comparison on the OVIS validation set. Bold and underline indicate the best and second-best performances, respectively. We use the online version for GenVIS [16] and DVIS [37].
[…]
Transformer Encoder 52.8 73.8 58.1 49.4 60.2
Backbone 54.5 77.2 60.3 50.9 61.7
OVIS.
Our method, VISAGE, is evaluated on the OVIS dataset. In Tab. 3, VISAGE demonstrates competitive performance when compared to other state-of-the-art methods [16, 23, 36]. It is important to note that these methods incorporate sequential information during training. Consequently, these methods shows marginally better performance than VISAGE, which is expected given that the OVIS dataset is characterized by its longer video sequences.
@@ 4.4 Ablation Studies
Appearance feature.
In Tab. 4, we analyze the impact of feature maps which create the appearance query. When we make the appearance query from the transformer encoder feature maps, there is a slight drop in performance. This indicates that backbone feature maps contain rich visual information compare to transformer encoder feature maps.
Appearance guidance.
Tab. 5 shows effectiveness of our appearance-guided enhancement. We omit appearance information by not computing the appearance similarity in line 9 of Algorithm 1, resulting in a matching process that relies solely on object similarity. As indicated by rows 1 and 3 in Tab. 5, the absence of appearance information leads to a degradation in performance. Further analysis of the effectiveness of appearance guidance is discussed in Sec. 4.5.
Memory bank.
In Tab. 5, we demonstrate the necessity of the memory bank for compensating for the lack of temporal information. We evaluate performance differences between scenarios with and without the use of the memory bank. Rows 1 and 2 in Tab. 5 show decreased performance compared to rows 3 and 4, respectively. If there is no memory bank, our method lacks access to historical information beyond the frame right before the current one, since it measures similarities solely between the current frame and the one immediately preceding it. By implementing a memory bank, our method gains awareness of previous frames, which leads to improved performance.
Appearance Memory Bank YouTube-VIS 2019
@@
54.5 77.2 60.3 50.9 61.7
Table 5: Ablation study on appearance guidance and memory bank utilization in YTVIS 2019 validation Set, with memory bank window size $W=5$.
@@
[29, 36] (w/o NMS) 52.7 75.3 57.6 49.3 61.6
[29, 36] (w/ NMS) 54.4 77.9 59.0 49.9 62.8
ours (w/o NMS) 54.5 77.2 60.3 50.9 61.7
Table 6: Ablation study of the tracker type on YTVIS 2019 validation set.
Simplified tracker.
Previous query-matching based methods [29, 36], extend the tracker from [8], which employs heuristic post-processing techniques such as NMS. In contrast, our tracker relies solely on a cosine similarity metric and the Hungarian algorithm for matching. To analyze the impact of such handcrafted design, we modify our inference pipeline to align with the approaches used in previous methods [29, 36]. As demonstrated in Tab. 6, both the previous strategies and our simplified version show comparable performance. However, removing the heuristic design element NMS from the previous approach results in performance degradation. This outcome implies that previous approaches rely on handcrafted design, unlike our simpler tracker.
@@ 4.5 Analysis of Appearance-Guided Enhancement
T-SNE Visualization.
As illustrated in Fig. 4, we analyze the impact of appearance-guided enhancement in VISAGE on its association capabilities using the OVIS dataset by visualizing the query embeddings for an entire video, with consistent colors indicating the same instance. The visualization includes three types of query embeddings: object-only, appearance-only, and a combination of both object and appearance queries. The embeddings that combine object and appearance query result in a well-clustered visualization. While the appearance-only query embeddings exhibit reasonable clustering, the integration of both types of queries leads to more distinct and pronounced clustering. This enhanced clustering indicates that our appearance-guided enhancement leads to better association.
@@
Track GenVIS [16] 54.8 70.6 58.8 61.2 65.9
CTVIS [36] 63.7 78.7 69.7 70.3 74.8
VISAGE (w/o app) 65.2 80.9 71.3 70.2 75.2
VISAGE 65.8 81.1 71.7 70.8 76.3
Swap GenVIS [16] 41.8 59.7 44.0 50.8 57.0
CTVIS [36] 53.1 71.7 56.9 61.7 65.7
VISAGE (w/o app) 51.8 72.1 54.4 57.7 62.8
VISAGE 63.0 79.5 69.3 67.8 73.5
Table 7: Comparisons on Pseudo dataset. Bold denote the highest accuracy.
@@
4.6 Qualitative Results
Fig. 6 displays the visualization results of our VISAGE across various challenging scenarios. In the first row, several cats are positioned closely together and the red-colored cat crosses another cat. In such cases, relying solely on location information can often result in identity-switching errors. However, our VISAGE tracks each cat accurately despite their adjacent positions and sequential intersections. In the second row, we present a similar scenario involving different classes: a person and a cow.
@@ 5 Limitations and Future Works
VISAGE addresses the insufficiency of existing query-based online VIS methods, specifically their lack of appearance awareness. However, our approach has two limitations. First, it assumes equal importance for appearance and location information for its simplicity. In some scenarios, appearance may be more critical than location, while in others, location may be prioritized. Designing a method to adaptively weight the importance of appearance and location information for stable tracking will be a future direction. Second, VISAGE focuses exclusively on the query-matching approach, and we observe that query-propagation methods also tend to rely on location information. Another future direction involves incorporating appearance information awareness into query-propagation approaches.

### 08 Mar 2024 v2

@@ 4.3 Main Results
We compare our methods with state-of-the-art methods on four VIS benchmarks: YTVIS 2019/2021/2022 and OVIS. The results are reported in Tab. 1 and Tab. 2.
4.3.1 Youtube-VIS 2019 & 2021 & 2022.
As shown in Tab. 1, we compare our VISAGE with previous state-of-the-art methods. On the YTVIS 2019 benchmark, VISAGE performs on par with the highest-performing method cited as CTVIS [37]. Notably, both VISAGE and CTVIS [37] surpass previous methods by a large margin. On the YTVIS 2021 benchmark, an improved version of YTVIS 2019, VISAGE outperforms other existing methods. We achieve the highest performance on both the YTVIS 2019 and 2021 datasets by incorporating appearance-guided enhancement. Finally, on the YTVIS 2022 benchmark, our VISAGE exhibits comparable performance compared to GenVIS[15] and TCOVIS [23].
4.3.2 OVIS.
In Tab. 2, we present a comparison on the OVIS benchmark [25], which is characterized by long videos and complex scenarios, including frequent occlusions. VISAGE achieves performance on par with the previously established state-of-the-art methods such as GenVIS [15], TCOVIS [23], and CTVIS [37]. Through our proposed properties, which include appearance-guided enhancement and a simplified tracker, VISAGE effectively handles such long and complicated videos. As a result, these advancements enable VISAGE to achieve state-of-the-art performance.
Table 2: Comparison on the OVIS validation set. Bold indicates the best performance. We use the online version for GenVIS [15] and DVIS [38].
Method OVIS
@@
VISAGE 36.2 60.3 35.3 17.0 40.3
@@ 4.4 Ablation Studies
4.4.1 Appearance feature.
In Tab. 3, we analyze the impact of feature maps which create the appearance query. When we make the appearance query from the transformer encoder feature maps, there is a degradation in performance. This indicates that backbone feature maps contain rich visual information compared to transformer encoder feature maps.
4.4.2 Appearance guidance.
Tab. 5 shows effectiveness of our appearance-guided enhancement. We omit appearance information by setting the $\alpha$ in line 9 of Algorithm 1 to 0, resulting in a matching process that relies solely on object similarity. As indicated by rows 1 and 3 in Tab. 5, the absence of appearance information leads to a degradation in performance. Further analysis of the effectiveness of appearance guidance is discussed in Sec. 4.5.
4.4.3 Memory bank.
In Tab. 5, we demonstrate the necessity of the memory bank for compensating for the lack of temporal information. We evaluate performance differences between scenarios with and without the use of the memory bank. Rows 1 and 2 in Tab. 5 show decreased performance compared to rows 3 and 4, respectively. Without a memory bank, our method is constrained to using only the immediately preceding frame for historical context, relying on similarities measured exclusively between consecutive frames. However, the introduction of a memory bank expands this capability by leveraging a broader historical perspective. Furthermore, the adoption of a memory bank significantly boosts the effectiveness of appearance information. This is because historical appearance information provides a more reliable basis for matching than only considering the appearance from the immediate preceding frame to recognize an identical object. By implementing a memory bank, our method gains awareness of previous frames, which leads to improved performance.
[…]
Transformer Encoder 51.4 72.4 56.7 49.8 60.7
Backbone 55.1 78.1 60.6 51.0 62.3
Table 4: Ablation study on appearance guidance and memory bank utilization, with memory bank window size $W=5$.
Table 5: Ablation study of the tracklet post-processing.
App Memory Bank YouTube-VIS 2019
@@
55.1 78.1 60.6 51.0 62.3
@@
[31, 37] 52.7 75.3 57.6 49.3 61.6
[31, 37] 54.4 77.9 59.0 49.9 62.8
ours 55.1 78.1 60.6 51.0 62.3
Table 5: Ablation study of the tracklet post-processing.
4.4.4 Tracklet processing.
Previous query-matching based methods [31, 37], build upon the tracker framework from [8], incorporating handcrafted thresholding and heuristic post-processing techniques, including Non-Maximum Suppression (NMS). In contrast, our tracker exclusively utilizes cosine similarity and the Hungarian algorithm for matching, as detailed in Algorithm 1. To understand the impact of handcrafted designs, we align our inference pipeline with those of methods previously used  [31, 37]. The results, as detailed in Tab. 5, reveal that our streamlined approach consistently surpasses the performance of the conventional tracker. A detailed examination, especially between rows 1 and 2 of Tab. 5, reveals that the absence of heuristic elements, such as NMS, leads to a decline in performance. However, row 3 reveals that our simplified tracker performs impressively even without heuristic design elements. This suggests that a simplified tracker is capable of achieving commendable performance without the complexity of heuristic design elements.
@@ 4.5 Analysis of Appearance-Guided Enhancement
4.5.1 Appearance Weight.
In our methodology, as detailed in line 9 of Algorithm 1, the hyperparameter $\alpha$ plays a crucial role by representing the weighting of the appearance similarity. By tuning the value of $\alpha$, we modulate the emphasis placed on the appearance information. In Tab. 5, we demonstrate how our appearance guidance significantly enhances performance on the YTVIS 2019 dataset. In this section, we further analysis the impact of appearance cue on robust matching by exploring various values of $\alpha$.
Tab. 6 highlights the critical role of both appearance and location information in tracking. We analyze the impact of appearance weight $\alpha$ on both the YTVIS 2019 and OVIS validation sets. There are two extreme cases to consider: relying solely on location information or exclusively on appearance information. We observe a reduced performance on both datasets when the $\alpha$ value is set to 0, with this reduction being more pronounced on the OVIS dataset. Given OVIS’s complex scenarios, such as frequent occlusions, the dataset’s intricacies make the contribution of appearance information especially significant. Conversely, setting $\alpha$ to 1, and thus relying only on appearance information, results in a significant drop in performance. Without the positional cue, the model only depends on appearance information for matching across frames, leading to potential ambiguities in making object tracklets.
Table 6: Analysis of Appearance Weight $\alpha$. The plot demonstrates the relationship between different appearance weight $\alpha$ settings and their corresponding AP scores on the YouTube-VIS 2019 and OVIS validation sets.
$\alpha$ YouTube-VIS 2019 OVIS
AP
[…]
AP
[…]
0.00
53.4
76.8
58.7
49.8
61.2
32.2
55.6
30.9
16.1
36.3
0.25
53.6
77.0
59.4
50.0
61.1
34.5
59.2
32.5
16.7
38.8
0.50
54.5
77.2
60.3
50.9
61.7
34.8
59.8
35.3
16.2
39.4
0.75
55.1
78.1
60.6
51.0
62.3
36.2
60.3
35.3
17.0
40.3
1.00
24.9
36.7
28.7
40.9
53.7
11.4
22.7
10.2
10.3
24.9
However, integrating both appearance and location information consistently surpasses these two extreme cases, highlighting the complementary strengths of using both cues in establishing robust tracklets. Notably, increasing the value of $\alpha$, and thereby the emphasis on appearance information, correlates with performance enhancements.
4.5.2 T-SNE Visualization.
As illustrated in Fig. 5, we analyze the impact of appearance-guided enhancement in VISAGE on its association capabilities using the OVIS dataset by visualizing the query embeddings for an entire video, with consistent colors indicating the same instance. The visualization includes three types of query embeddings: object-only, appearance-only, and a combination of both object and appearance queries. While object-only embeddings for the same object exhibit scattered and inconsistent clustering, appearance-only query embeddings demonstrate a high degree of clustering. The integration of object and appearance query embeddings results in even more distinct and pronounced clustering, as exemplified by the red circle in Fig. 5. This enhanced clustering clearly indicates that our appearance-guided enhancement leads to better association.
4.5.3 Pseudo Dataset.
@@
[…]
Track
GenVIS [15]
54.8
70.6
58.8
61.2
65.9
CTVIS [37]
63.7
78.7
69.7
70.3
74.8
VISAGE
65.2
80.9
71.3
70.2
75.2
VISAGE
65.8 81.1 71.7 70.8 76.3
Swap
GenVIS [15]
41.8
59.7
44.0
50.8
57.0
CTVIS [37]
53.1
71.7
56.9
61.7
65.7
VISAGE
51.8
72.1
54.4
57.7
62.8
VISAGE
63.0 79.5 69.3 67.8 73.5
4.6 Qualitative Results
Fig. 6 displays the visualization results of our VISAGE across various challenging scenarios. In the first row, several cats are positioned closely together and the red-colored cat crosses another cat. In such cases, relying solely on location information can often result in identity-switching errors. In the second row, we present a similar scenario involving different classes: a person and a cow.
@@ 5 Limitations and Future Works
VISAGE addresses the lack of appearance awareness in existing query-based online VIS methods. Nevertheless, our approach encounters two main limitations. Firstly, while VISAGE primarily employs a query-matching strategy, we recognize that query-propagation methods similarly depend on location information. This observation points towards a future research direction: integrating appearance information awareness into query-propagation approaches. Secondly, a fundamental issue not unique to our method but prevalent in tracking-by-detection approaches, especially in recent query-based VIS methods, is the heavy reliance on frame-level detectors. During the tracking stage, these methods do not account for their own errors, allowing any inaccuracies to adversely affect video-level predictions. Addressing this challenge requires thorough consideration and represents another avenue for future work.